OpenStudy (anonymous):

11^9 / 11^ 8 simplify using the rules for exponets

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{11^9}{11^8} =11^{9-8}=? }\)

OpenStudy (solomonzelman):

finish it from here.